# PuLID: Pure and Lightning ID Customization via Contrastive Alignment

### v1

@@ Abstract
We propose Pure and Lightning ID customization (PuLID), a novel tuning-free ID customization method for text-to-image generation. By incorporating a Lightning T2I branch with a standard diffusion one, PuLID introduces both contrastive alignment loss and accurate ID loss, minimizing disruption to the original model and ensuring high ID fidelity. Experiments show that PuLID achieves superior performance in both ID fidelity and editability. Another attractive property of PuLID is that the image elements (e.g., background, lighting, composition, and style) before and after the ID insertion are kept as consistent as possible. Codes and models will be available at https://github.com/ToTheBeginning/PuLID.
@@ 2 Related Work
Tuning-based Text-to-image ID Customization. ID Customization for text-to-image models aims to empower pre-trained models to generate images of specific identities while following the text descriptions. Two seminal tuning-based works [5, 30] strive towards this goal. Textual Inversion [5] optimizes a new word embedding for the user-provided ID, and Dreambooth [30] fine-tunes the entire generator to further enhance fidelity. Subsequently, various approaches [12, 17, 8, 35] have explored different fine-tuning paradigms in the generator and embedding space to achieve superior ID fidelity and text alignment. Despite these advancements, the time-consuming optimization process for each ID, taking at least several minutes, restricts its broader application.
Tuning-free Text-to-image ID Customization. To ease the resource demand necessitated by online tuning, a series of tuning-free methods [36, 38, 25, 42, 20, 41, 3] have emerged, which directly encode ID information into the generation process. The major challenge these methods encounter is minimizing disruption to the original behavior of T2I models while still maintaining high ID fidelity.
@@ 3.4 Uncontaminated ID Insertion via Contrastive Alignment
Figure 3: Effect of $\mathcal{L}_{\text{align-sem}}$ and $\mathcal{L}_{\text{align-layout}}$.
Our alignment loss consists of two components: the semantic alignment loss and the layout alignment loss. We use textual features $K$ to query the UNET features $Q$. For each token in $K$, it will calculate the correlation with $Q$, and further aggregate $Q$ based on the correlation matrix. Analogous to Eq. 2, the attention mechanism here can be expressed as $\text{Attention}(K,Q,Q)$, which can be interpreted as the response of the UNET features to the prompt. The insight behind our semantic alignment loss is simple: if the embedding of ID does not affect the original model’s behavior, then the response of the UNET features to the prompt should be similar in both paths. Therefore, our semantic alignment loss $\mathcal{L}_{\text{align-sem}}$ can be formulated as follows:
@@
As illustrated in Fig. 3, the introduction of $\mathcal{L}_{\text{align-sem}}$ significantly mitigates the issue of ID information contaminating the model’s behavior. However, it cannot guarantee layout consistency, so we add a layout alignment loss $\mathcal{L}_{\text{align-layout}}$, which is defined as:
@@ 3.5 Optimizing ID Loss in a More Accurate Setting
$\mathcal{L}_{\text{id}}=CosSim\left(\phi(C_{id}),\phi(\text{L-T2I}(x_{T},C_{id},C_{txt}))\right),$ (6)
@@ 3.6 Full Objective
$\mathcal{L}_{\text{id}}=\mathcal{L}_{\text{diff}}+\mathcal{L}_{\text{align}}+\lambda_{\text{id}}\mathcal{L}_{\text{id}}.$ (7)
@@ 4.2 Test Settings
For consistency in comparison, unless otherwise specified, all the results in this paper are generated with the SDXL-Lightning [21] base model over $4$ steps using the DPM++ 2M sampler [15]. The CFG-scale is set to $1.2$, as recommended by [21]. Moreover, for each comparison sample, all methods utilize the same seed. We find that the comparison methods, namely InstantID [38] and IPAdapter (more specifically, IPAdapter-FaceID [1]) are highly compatible with the SDXL-Lightning model. Compared to using SDXL-base [26] as the base model, employing SDXL-Lightning results in InstantID generating more natural and aesthetically pleasing images, and enables IPAdapter to achieve higher ID fidelity. We also provide a quantitative comparison with these methods on SDXL-base, and the conclusions remaining consistent with those on SDXL-Lightning.
@@ 4.3 Qualitative Comparison
As shown in Fig. 4, when compared to SOTA methods such as IPAdapter and InstantID, our PuLID tends to achieve higher ID fidelity while creating less disruption to the original model. From rows 1, 2, 5, 6, and 7, it is clear that our method can attain high ID similarity in realistic portrait scenes and delivers better aesthetics. Conversely, other methods either fall short in ID fidelity or show diminished aesthetics compared to the base model. Another distinct advantage of our approach is that as the disruption to the model decreases, the results produced by PuLID accurately replicate the lighting (1st row), style (4th row), and even layout (5th row) of the original model. In contrast, although comparative methods can also perform stylization, notable style degradation can be noticed when compared to the original model. Finally, our model also possesses respectable prompt-editing capabilities, such as changing orientation (2nd row), altering attributes (6th row), and modifying accessories (7th row). More qualitative comparison can be found in Fig. 1.
@@ 4.4 Quantitative Comparison
Table 1: Quantitative comparisons. Comparison of ID cosine similarity with SOTA methods across different base models and datasets. PuLID (maximum ID sim) represents the model from second training stage.
SDXL-Lightning SDXL-base
DivID-120 Unsplash-50 DivID-120 Unsplash-50
PhotoMaker - - 0.271 0.193
IPAdapter 0.619 0.615 0.597 0.572
InstantID 0.725 0.614 0.755 0.648
PuLID (maximum ID sim) 0.761 0.708 0.773 0.711
PuLID (ours) 0.733 0.659 0.734 0.666
To quantitatively compare the methods, we adopt the ID cosine similarity to evaluate ID fidelity, with ID embeddings extracted using CurricularFace [13]. CurricularFace is different from the face recognition models we use in the ID encoder and for ID loss calculation. Table 1 presents the quantitative results. As seen in the table, our model that strives for maximum ID fidelity (the model from the second training stage) outperforms existing methods on all test sets and base models. Even after introducing the alignment loss and sacrificing some ID fidelity, our final model, PuLID, still outperforms the comparative methods in most scenarios, except for when employing SDXL-base as the base model and being slightly inferior to InstantID on DivID-120.
Moreover, Table 1 shows that InstantID and IPAdapter are well compatible with SDXL-Lightning. IPAdapter exhibits some metric improvements after utilizing SDXL-Lightning. Despite a minor decline in metrics when transitioning to SDXL-Lightning, InstantID sees substantial improvements in image quality and usability (refer to supp for more details). Furthermore, our method still outperforms InstantID when using SDXL-base as the base model. We observed that PhotoMaker shows limited compatibility with SDXL-Lightning, suffering significant performance degradation as a result. Hence, we only compare its performance on SDXL-base in this table.
4.5 More Applications
Figure 5: More applications. Including style changes, IP fusion, accessory modification, recontextualization, attribute editing, transformation from non-photo-realistic domain to photo-realistic domain, and ID mixing. Note that all these high-quality images are generated in just $4$ steps with SDXL-Lightning model, without the need for additional Lora.
We provide more applications of our PuLID in Fig. 5, encompassing style alterations (1st row), IP fusion (2nd row), accessories modification (3rd row), recontextualization (4th row), attributes editing (5th row), transformation from non-photo-realistic domains to photo-realistic ones (6th row), and ID mixing (7th row).
@@ 4.6 Ablation
Alignment loss ablation. Fig. 6 displays a comparison between models trained with and without the alignment loss $\mathcal{L}_{\text{align}}$. As observed, without $\mathcal{L}_{\text{align}}$, the embedding of ID severely disrupts the behavior of the original model. This disruption manifests as an inability for the prompt to precisely modify style (columns 2-3) and orientation (column 4). Also, the layout would collapse to the extent that the face occupies the majority of the image area, resulting in a diminished diversification of the layout. However, with the introduction of our alignment loss, this disruption can be significantly reduced.
ID loss ablation. Table 2 illustrates the improvement in ID fidelity using the naive ID loss (directly predicting $x_{0}$ from current timestep) and the more accurate ID loss $\mathcal{L}_{\text{id}}$ introduced in this paper, in comparison to the baseline. As observed, $\mathcal{L}_{\text{id}}$ can accomplish a greater improvement compared to the naive ID loss. We attribute this to the more precise $x_{0}$ provided by the Lightning-T2I branch, which also better aligns with the testing setting, thereby making the optimization of ID loss more direct and effective.
Figure 6: Alignment loss ablation.
Table 2: ID loss ablation.
@@
Baseline (Stage1) 0.561 0.514
w/ ID Loss naive 0.652 0.601
w/ ID Loss (Stage2) 0.761 0.708

### v2

@@ Abstract
We propose Pure and Lightning ID customization (PuLID), a novel tuning-free ID customization method for text-to-image generation. By incorporating a Lightning T2I branch with a standard diffusion one, PuLID introduces both contrastive alignment loss and accurate ID loss, minimizing disruption to the original model and ensuring high ID fidelity. Experiments show that PuLID achieves superior performance in both ID fidelity and editability. Another attractive property of PuLID is that the image elements (e.g., background, lighting, composition, and style) before and after the ID insertion are kept as consistent as possible. Codes and models are available at https://github.com/ToTheBeginning/PuLID.
@@ 2 Related Work
Tuning-free Text-to-image ID Customization. ID Customization for text-to-image models aims to empower pre-trained models to generate images of specific identities while following the text descriptions. To ease the resource demand necessitated by tuning-based methods [7, 36, 14, 19, 10, 41], a series of tuning-free methods [42, 44, 29, 50, 22, 48, 51, 4] have emerged, which directly encode ID information into the generation process. The major challenge these methods encounter is minimizing disruption to the original behavior of T2I models while still maintaining high ID fidelity.
@@ 3.4 Uncontaminated ID Insertion via Contrastive Alignment
Figure 3: Illustration and Effect of the alignment loss.
Our alignment loss consists of two components: the semantic alignment loss and the layout alignment loss. An illustration is presented in Fig. 3 (a). We use textual features $K$ to query the UNET features $Q$. For each token in $K$, it will calculate the correlation with $Q$, and further aggregate $Q$ based on the correlation matrix. Analogous to Eq. 2, the attention mechanism here can be expressed as $\text{Attention}(K,Q,Q)$, which can be interpreted as the response of the UNET features to the prompt. The insight behind our semantic alignment loss is simple: if the embedding of ID does not affect the original model’s behavior, then the response of the UNET features to the prompt should be similar in both paths. Therefore, our semantic alignment loss $\mathcal{L}_{\text{align-sem}}$ can be formulated as follows:
@@
As illustrated in Fig. 3 (b), the introduction of $\mathcal{L}_{\text{align-sem}}$ significantly mitigates the issue of ID information contaminating the model’s behavior. However, it cannot guarantee layout consistency, so we add a layout alignment loss $\mathcal{L}_{\text{align-layout}}$, which is defined as:
@@ 3.5 Optimizing ID Loss in a More Accurate Setting
$\mathcal{L}_{\text{id}}=1-CosSim\left(\phi(C_{id}),\phi(\text{L-T2I}(x_{T},C_{id},C_{txt}))\right),$ (6)
@@ 3.6 Full Objective
$\mathcal{L}=\mathcal{L}_{\text{diff}}+\mathcal{L}_{\text{align}}+\lambda_{\text{id}}\mathcal{L}_{\text{id}}.$ (7)
@@ 4.2 Test Settings
For consistency in comparison, unless otherwise specified, all the results in this paper are generated with the SDXL-Lightning [23] base model over $4$ steps using the DPM++ 2M sampler [17]. The CFG-scale is set to $1.2$, as recommended by [23]. Moreover, for each comparison sample, all methods utilize the same seed. We observe that the comparison methods, namely InstantID [44] and IPAdapter (more specifically, IPAdapter-FaceID [1]) are highly compatible with the SDXL-Lightning model. As shown in appendix subsection 7.5, when compared to using SDXL-base [30] as the base model, employing SDXL-Lightning results in InstantID generating more natural and aesthetically pleasing images, and enables IPAdapter to achieve higher ID fidelity. Furthermore, we provide a quantitative comparison with these methods on SDXL-base, with the conclusions remaining consistent with those on SDXL-Lightning.
@@ 4.3 Qualitative Comparison
As shown in Fig. 4, when compared to SOTA methods such as IPAdapter and InstantID, our PuLID tends to achieve higher ID fidelity while creating less disruption to the original model. From columns 1, 2, 5, 6, and 7, it is clear that our method can attain high ID similarity in realistic portrait scenes and delivers better aesthetics. Conversely, other methods either fall short in ID fidelity or show diminished aesthetics compared to the base model. Another distinct advantage of our approach is that as the disruption to the model decreases, the results produced by PuLID accurately replicate the lighting (1st column), style (4th column), and even layout (5th column) of the original model. In contrast, although comparative methods can also perform stylization, notable style degradation can be noticed when compared to the original model. Finally, our model also possesses respectable prompt-editing capabilities, such as changing orientation (2nd column), altering attributes (6th column), and modifying accessories (7th column). More qualitative results can be found in subsection 7.1 and subsection 7.2 of the appendix.
@@ 4.4 Quantitative Comparison
Table 1: Quantitative comparisons. *We observed that PhotoMaker shows limited compatibility with SDXL-Lightning, hence, we compare its performance on SDXL-base in this table.
DivID-120 Unsplash-50
[…]
PhotoMaker* 0.271 26.06 0.649 0.193 27.38 0.692
IPAdapter 0.619 28.36 0.703 0.615 28.71 0.701
InstantID 0.725 28.72 0.680 0.614 30.55 0.736
PuLID (ours) 0.733 31.31 0.812 0.659 32.16 0.840
The quantitative results are reported in Table 1. Face Sim. represents the ID cosine similarity, with ID embeddings extracted using CurricularFace [15]. CurricularFace is different from the face recognition models we use in the ID encoder and for calculating ID loss. CLIP-T [32] measures the ability to follow prompt. We also use CLIP-I to quantify the CLIP image similarity between two images before and after the ID insertion. A higher CLIP-I metric indicates a smaller modification in image elements (such as the background, composition, style) after ID insertion, suggesting a lower degree of disruption to the original model’s behavior. We provide the evaluation prompts in the appendix. As observed from Table 1, our method, PuLID, surpasses comparison methods across all three metrics, achieves SOTA performance in terms of both ID fidelity and editability. Furthermore, our method significantly outperforms others with respect to the CLIP-I metric, implying that our method incurs much less intrusion on model behavior compared to other methods.
@@ 4.5 Ablation
Alignment loss ablation. Fig. 5 displays a qualitative comparison between models trained with and without the alignment loss $\mathcal{L}_{\text{align}}$. As observed, without $\mathcal{L}_{\text{align}}$, the embedding of ID severely disrupts the behavior of the original model. This disruption manifests as an inability for the prompt to precisely modify style (the left two cases of Fig. 5) and orientation (the lower right case of Fig. 5). Also, the layout would collapse to the extent that the face occupies the majority of the image area, resulting in a diminished diversification of the layout. However, with the introduction of our alignment loss, this disruption can be significantly reduced. From Table 2, we could also observe a large improvement in CLIP-T and CLIP-I when equipped with $\mathcal{L}_{\text{align}}$ (from Stage2 to Stage3).
ID loss ablation. Table 2 illustrates the improvement in ID fidelity using the naive ID loss (directly predicting $x_{0}$ from current timestep) and the more accurate ID loss $\mathcal{L}_{\text{id}}$ introduced in this paper, in comparison to the baseline. As observed, $\mathcal{L}_{\text{id}}$ can accomplish a greater improvement compared to the naive ID loss. We attribute this to the more precise $x_{0}$ provided by the Lightning-T2I branch, which also better aligns with the testing setting, thereby making the optimization of ID loss more direct and effective. Another worth mentioning point is that the baseline, naively trained on the internal dataset, underperforms in ID fidelity and editability. Conversely, the introduction of the PuLID training paradigms delivers significant enhancement. Therefore, this substantiates that the improvement mainly comes from the method, rather than the dataset.
Figure 5: Qualitative comparison for ablation study on alignment loss.
Table 2: Quantitative comparisons for ablation studies on ID loss and alignment loss.
@@
[…]
Baseline (Stage1) 0.561 29.06 0.736 0.514 30.16 0.769
w/ ID Loss naive 0.652 27.05 0.683 0.601 28.00 0.707
w/ ID Loss (Stage2) 0.761 24.91 0.624 0.708 25.83 0.646
PuLID (Stage3) 0.733 31.31 0.812 0.659 32.16 0.840
5 Limitation
While our PuLID achieves superior ID fidelity and editability with minimal disruption to the base model’s behavior, it still has limitations. Due to the incorporation of the Lightning T2I branch, the training speed per iteration is slower than that of conventional diffusion training, and it also demands more CUDA memory. However, this issue can be significantly mitigated when future fast sampling methods can generate satisfying results in a single step (currently, we use 4 steps). Another limitation is that although the accurate ID loss introduced in the Lightning T2I branch markedly enhances ID fidelity, it also impacts image quality to some extent, such as causing blurriness in faces. This can be discerned in Fig. 3 (b) and Fig. 5. Nonetheless, this issue is not uniquely associated to our method, as similar phenomena has been observed in reward tuning methods [49, 31, 5, 47], with [5] attributing it to the reward-hacking problem. Although this issue can be largely alleviated by pairing with our proposed contrastive alignment loss, future work could explore a more effective ID loss that does not negatively affect image quality.
6 Conclusion
This paper presents PuLID, a novel tuning-free approach to ID customization in text-to-image generation. By incorporating a Lightning T2I branch along with a contrastive alignment strategy, PuLID achieves superior ID fidelity and editability with minimal disruption to the base model’s behavior. Experimental results demonstrate that PuLID surpasses current methods, showcasing its potential for flexible and efficient personalized image generation. Future work could explore the application of this proposed training paradigm to other image customization tasks, like IP and style customization.